CBA Bank is studying the use of automatic teller machines. Of interest is whether young adults (under 25 yrs) use the machine more than senior. To investigate further, sample of customers under 25 yrs of age an customers over 60 yrs of age were selected.
The number of ATM transactions last month was determined for each selected individual and the results are shown below.

Age Number of transactions
Under 25 10 10 11 15 7 11 10 9
Over 60 4 4 7 7 4 5 1 7 10 5

At the 0.01 significant level. Can bank management conclude that younger customers use the ATM more?

Answer by stanbon(48568)

To investigate further, sample of customers under 25 yrs of age an customers over 60 yrs of age were selected.
The number of ATM transactions last month was determined for each selected individual and the results are shown below.
Age Number of transactions
Under 25 10 10 11 15 7 11 10 9
Over 60 4 4 7 7 4 5 1 7 10 5
At the 0.01 significant level. Can bank management conclude that younger customers use the ATM more?
---------------------
Ho: mu(under) - mu(over)=0
Ha: mu(under)-mu(over)>0
mu(under)=12 ; mu(over)=10.36
z(12-10.36)=z(1.64)=0.4104457...
p value= 0.34
---------
Fail to Reject Ho
Use by younger and older is the same
----------------
Cheers,
Stan H.